# compat=1.0.0

### v1

@@
Improved constraint on Higgs boson self-couplings with quartic and cubic power dependence in the cross section
Hai Tao Li1, Zong-Guo Si1, Jian Wang1,2, Xiao Zhang1, Dan Zhao1
1School of Physics, Shandong University, Jinan, Shandong 250100, China
2Center for High Energy Physics, Peking University, Beijing 100871, China
Precise information on the Higgs boson self-couplings provides the foundation for unveiling the electroweak symmetry breaking mechanism. Due to the scarcity of Higgs boson pair events at the LHC, only loose limits have been obtained. This is based on the assumption that the cross section is a quadratic function of the trilinear Higgs self-coupling in the $\kappa$ framework. However, if higher-order corrections of virtual Higgs bosons are included, the function form would dramatically change. In particular, new quartic and cubic power dependence on the trilinear Higgs self-coupling would appear. To get this new function form, we have performed a specialized renormalization procedure suitable for tracking all the Higgs self-couplings in each calculation step. Moreover, we introduce renormalization of the scaling parameter in the $\kappa$ framework to ensure the cancellation of all ultraviolet divergences. With the new function forms of the cross sections in both the gluon-gluon fusion and vector boson fusion channels, the upper limit of $\kappa_{\lambda_{3}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ by the ATLAS (CMS) collaboration is reduced from 6.6 (6.49) to 5.4 (5.37). However, it is still hard to extract a meaningful constraint on the quartic Higgs self-coupling $\lambda_{\rm 4H}$ from Higgs boson pair production data. We also present the invariant mass distributions of the Higgs boson pair at different values of $\kappa_{\lambda}$, which could help to set optimal cuts in the experimental analysis.
@@ 1 Introduction
After the discovery of the Higgs boson at the Large Hadron Collider (LHC) [1, 2], precise measurements of its properties, such as the mass, spin, and couplings to gauge bosons and fermions, have become critically important [3, 4, 5, 6, 7, 8, 9]. To date, these measurements have been consistent with the expectations of the Standard Model (SM) [10, 11]. However, the trilinear and quartic Higgs self-couplings, i.e., $\lambda_{\rm 3H}$ and $\lambda_{\rm 4H}$, which play special roles in the spontaneous electroweak (EW) symmetry breaking, stability of the universe [12] and precision test of the SM, are still subject to large uncertainties.
@@
A lot of efforts have been devoted to improving the measurement of the Higgs self-coupling. The most direct approach involves measuring the cross section of Higgs boson pair production. The dominant channel is the production via gluon-gluon fusion (ggF). At the leading order (LO), the process occurs via a top-quark loop. In the large $m_{t}$ limit, using so called Higgs effective field theory, the cross section of ggF Higgs boson pair production is known up to next-to-next-to-next-to-leading-order (N3LO) QCD corrections  [13, 14, 15, 16, 17]. The soft gluon resummation effect was also calculated [18, 19, 20]. With full top quark mass dependence, only next-to-leading-order (NLO) QCD correction is currently available [21, 22, 23, 24]. Estimates of the top-quark mass effects at next-to-next-to-leading-order (NNLO) have been performed by [25, 26, 27, 28, 29]. To provide an accurate simulation of the events, a fully differential calculation of the Higgs boson pair production and decay to the $b\bar{b}\gamma\gamma$ final state with exact top quark mass at QCD NLO was presented [30]. In addition, the NLO EW corrections have been investigated in [31, 32, 33, 34, 35, 36, 37]. The subdominant channel—vector boson fusion (VBF)—has also been calculated up to N3LO in QCD [38, 39, 40, 41].
@@
Using the data of Higgs boson pair production at the Run 2 LHC, the current best constraints are $-0.6<\kappa_{\lambda_{\rm 3H}}<6.6$ [42] and $-1.24<\kappa_{\lambda_{\rm 3H}}<6.49$ [11] extracted from the ATLAS and CMS measurements, respectively, in the $\kappa$ framework, in which $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ with $\lambda_{\rm 3H}^{\rm SM}$ being the SM value of the trilinear Higgs self-coupling. On the one hand, the process of single Higgs boson production and decay depends on the Higgs self-coupling via higher-order EW corrections and can also provide certain constraint [43, 44, 45, 46, 47, 48]. A measurement performed by CMS using the differential fiducial cross section in bins of $p_{T}^{H}$ sets the constraint as $-5.4<\kappa_{\lambda_{\rm 3H}}<14.9$ [49]. A combined analysis of the single Higgs and double Higgs data by ATLAS shows that the constraint on the Higgs self-coupling is $-0.4<\kappa_{\lambda_{\rm 3H}}<6.3$ at $95\%$ confidence level with the assumption that the new physics changes only the Higgs self-coupling [42].
@@
The above constraints are obtained based on the fact that the cross section of Higgs boson pair production can be taken as a function of the Higgs self-couplings. Indeed, even with higher-order QCD corrections, the cross section is a quadratic function of the trilinear Higgs self-coupling $\lambda_{\rm 3H}$. Nevertheless, higher-order EW corrections incorporate contributions from Feynman diagrams containing one or more triple Higgs or quadruple Higgs vertices. These corrections exhibit a distinct functional dependence on the Higgs self-coupling. Specifically, new quartic and cubic dependence on the trilinear Higgs self-coupling would appear, which has a significant impact on the constraints given that the present upper limit is so large.
However, in practice, it is not obvious to keep the explicit dependence on the Higgs self-coupling, which is considered a derived parameter in the conventional calculation of EW corrections, especially in the renormalization procedure; see, e.g., Refs. [50, 35]. In the SM, the Higgs boson self-coupling $\lambda$ is determined by the Higgs boson mass and the vacuum expectation value. Therefore, the calculation in the SM can not be simply extended to the case with general $\lambda_{\rm 3H}$ and $\lambda_{\rm 4H}$, and the result with higher power (beyond quadratic) dependence on the Higgs self-couplings in the general $\kappa$ framework is still unavailable. In this work, we propose a renormalization procedure that explicitly keeps the Higgs self-couplings, and calculate the cross sections of the Higgs boson pair production via both ggF and VBF. Our results show that the upper limit of $\kappa_{\lambda_{\rm 3H}}$, which is 6.6 by ATLAS (6.49 by CMS), can be reduced by about 20% after including higher power dependence of the cross section on the Higgs self-couplings.
@@ 2 Renormalization in the $\kappa$ framework
The LO contribution to the ggF Higgs boson pair production $g(p_{1})g(p_{2})\to H(p_{3})H(p_{4})$ comes from the top-quark induced triangle and box Feynman diagrams, which are of order $\lambda_{\rm 3H}$ and $\lambda_{\rm 3H}^{0}$, respectively. Consequently, the LO cross section at the 13 TeV LHC is given by
@@
The subscript $\lambda_{\rm 3H}$ in $\kappa_{\lambda_{\rm 3H}}$ indicates that this factor measures the deviation of the trilinear Higgs coupling. Below we will also need $\kappa_{\lambda_{\rm 4H}}$ to represent the modification of the quartic Higgs self-coupling. The SM cross section is reproduced at $\kappa_{\lambda_{\rm 3H}}=1$ and $\kappa_{\lambda_{\rm 4H}}=1$. The quadratic function form persists even if higher-order QCD corrections are included [51], e.g.,
$\sigma_{\rm ggF,NNLO-FT}^{\kappa_{\lambda}}=(~{}10.8~{}\kappa_{\lambda_{\rm 3H}}^{2}-49.6~{}\kappa_{\lambda_{\rm 3H}}+70.0~{})~{}~{}{\rm fb},$ (2)
where the full one-loop real contributions are combined with the other NNLO QCD corrections in the large $m_{t}$ limit. Compared with Eq. (1), the QCD corrections are so significant that each term in $\kappa_{\lambda_{\rm 3H}}$ has more than doubled.
The EW corrections would change the above function form in two aspects. First, the coefficients of the quadratic, linear, and constant terms would be changed. However, the magnitude is only a few percent as indicated by the calculation in Ref. [35]. This kind of correction marginally affects the constraints on Higgs self-coupling and thus is neglected in our study. Second, higher power or new dependence on the Higgs self-couplings arises. One can see in Fig. 1 some typical two-loop Feynman diagrams, which would make contributions of order $\lambda_{\rm 3H}^{3}$, $\lambda_{\rm 3H}^{2}$, $\lambda_{\rm 4H}\lambda_{\rm 3H}$, $\lambda_{\rm 4H}$ to the amplitudes. As a result, the cross section would then contain the quartic and cubic $\lambda_{\rm 3H}$ terms and start to be sensitive to the quartic Higgs self-coupling $\lambda_{\rm 4H}$. It is the purpose of our work to investigate these corrections, which are labeled by $\delta\sigma_{\rm EW}^{\kappa_{\lambda}}$ below, and their impact on the constraints for the Higgs self-couplings.
@@
Our calculation of the two-loop diagrams proceeds as follows. We use FeynArts [52] to generate the Feynman diagrams and amplitudes. The amplitudes are written as a linear combination of two tensor structures with the coefficients called form factors. After performing the Dirac algebra with FeynCalc [53, 54, 55], we are left with scalar integrals for each form factor. We did not attempt to reduce all of them to master integrals and construct the differential equations for the master integrals as in conventional methods. Instead, we chose to calculate the scalar integrals directly with the numerical package AMFlow [56, 57] for specific phase space points. The main reason is that the construction of the differential equation is time-consuming with general kinematic dependence. Even if the differential equation is obtained, the analytical solution seems not feasible with current technologies. Numerical solutions of the differential equations often suffer from accuracy loss. Direct numerical calculation at each phase space point ensures accuracy. The challenge lies in covering the entire phase space efficiently. Fortunately, the process of $gg\to HH$ is dominated by the S-wave component, rendering the amplitude insensitive to the scattering angle. Moreover, its dependence on the scattering energy is also weak, except in very high-energy regions, as illustrated below. These characteristics allow us to generate a grid with a limited data set, which can be used to accurately calculate the amplitude at any point in the phase space.
Notice that the sum of all one-particle irreducible two-loop diagrams is finite. But the one-particle reducible two-loop diagrams contain ultraviolet divergences. They will cancel after considering the contributions from the counter-terms in renormalization.
@@
where $Z_{v}$ is the renormalization constant for the vacuum expectation value. The renormalized Lagrangian in the $\kappa$ framework after EW gauge symmetry breaking is written as
$\mathcal{L}_{\rm H}$ $=\frac{1}{2}Z_{\phi}(\partial_{\mu}H)^{2}-\left(-\frac{1}{2}Z_{\mu^{2}}Z_{\phi}Z_{v}^{2}\mu^{2}v^{2}+\frac{1}{4}Z_{\lambda}Z_{\phi}^{2}Z_{v}^{4}\lambda v^{4}\right)-(Z_{\lambda}Z_{\phi}^{2}Z_{v}^{3}\lambda v^{3}-Z_{\mu^{2}}Z_{\phi}Z_{v}\mu^{2}v)H$
@@
where the ellipsis represents the terms involving EW gauge bosons. Note that the letter $\lambda$ is solely used for the Higgs self-coupling in the SM but $\lambda_{\rm 3H}\equiv\kappa_{\lambda_{3}}\lambda$ and $\lambda_{\rm 4H}\equiv\kappa_{\lambda_{4}}\lambda$ denote the Higgs self-couplings that could be modified by new physics. We have added renormalization constants $Z_{\kappa_{\rm 3H}}$ and $Z_{\kappa_{\rm 4H}}$ for the rescaling parameters $\kappa_{\lambda_{3}}$ and $\kappa_{\lambda_{4}}$ to account for potential new physics effect in renormalization. Following the general principle of the $\kappa$ framework, we have assumed that the new physics does not affect the vacuum expectation value and Higgs mass when we rescale the Higgs self-couplings. The second term in the first line does not contain any field and thus can be safely dropped. Writing $Z=1+\delta Z$, the third term can be expanded as
@@
where we have neglected higher-order corrections that are products of two $\delta Z$. We choose the renormalization condition such that there is no tadpole contribution. This condition requires $\mu^{2}=\lambda v^{2}$ at the tree level, and $(\delta Z_{\mu^{2}}-\delta Z_{\lambda}-\delta Z_{\phi}-2\delta Z_{v})\mu^{2}v+T=0$ at the one-loop level with $T$ being the contribution from the one-loop tadpole diagrams. $\delta Z_{v}$ would be determined after considering the renormalization of the EW gauge sector. Since we focus on the corrections induced by the Higgs self-coupling, we can simply take $\delta Z_{v}+\delta Z_{\phi}/2=0$. We adopt dimensional regularization, i.e., setting the space-time dimension $d=4-2\epsilon$, to symbolize the ultraviolet divergence and take $\mu_{R}$ as the renormalization scale. The tadpole diagram is evaluated to be
@@
$\frac{1}{2}(\partial_{\mu}H)^{2}-\mu^{2}H^{2}-\left(\frac{3}{2}\delta Z_{\lambda}+\frac{5}{2}\delta Z_{\phi}-\frac{1}{2}\delta Z_{\mu^{2}}+3\delta Z_{v}\right)\mu^{2}H^{2}$
@@
where $s=(p_{1}+p_{2})^{2},\beta=\sqrt{1-4m_{H}^{2}/s}$ and $C_{0}[m_{H}^{2},m_{H}^{2},s,m_{H}^{2},m_{H}^{2},m_{H}^{2}]$ is a scalar integral that can be calculated by Package-X [58]. $\mathcal{M}_{gg\to H^{*}\to HH}^{\rm LO}$ is the LO amplitude which contains the Higgs self-coupling. In the SM, the divergences in the first line vanish and $\delta Z_{\kappa_{\rm 3H}}$ is not needed. In the general $\kappa$ framework, it is essential to include $\delta Z_{\kappa_{\rm 3H}}$ in renormalization. We adopt the $\overline{\rm MS}$ scheme to subtract the divergences. As a result, the rescaling parameter is scale-dependent. If it is expressed in terms of the value at the scale $m_{H}$, then an additional contribution from its perturbative expansion cancels the above $\ln(\mu_{R}^{2}/m_{H}^{2})$ term exactly.
@@
Now it is ready to compute the finite part of the squared amplitudes. We set the two-dimensional grid as a function of the Higgs velocity $\beta$ and $\cos\theta$ with $\theta$ the scattering angle. The range of $\beta$ is from 0 to 1, and $\cos\theta$ is also taken from 0 to 1 since the squared amplitudes are symmetric under $\theta\to\pi-\theta$. The grids for the LO squared amplitudes and ${\lambda}$ dependent EW corrections are illustrated in Fig. 2. We see that the squared amplitudes are stable against the change within the region $\beta<0.6$. For larger $\beta$, the LO amplitudes rise dramatically and then start to drop when $\beta>0.9$. In contrast, the $\lambda$ dependent correction first decreases and then increases when $\beta$ is larger than 0.85. These variations are mainly due to the large logarithms $\ln^{i}(1-\beta)$. And therefore we have constructed the grid as a function of $\ln^{i}(1-\beta)$ for $\beta\geq 0.96$. We have tested the grid by comparing the generated values [^1] The Lagrange interpolation method is used to generate the squared amplitudes at the phase space point not on the lattice. and the ones obtained by direct high-precision computation for the LO squared amplitude at some phase space points that are not on the grid lattice. We find good agreement at the per-mille level. We have used the grid to calculate the total cross section by performing the convolution with the parton distribution function (PDF) and phase space integrations. Comparing it to the result obtained using analytical expressions or the OpenLoops package [59, 60, 61], we find that the relative difference is lower then $\mathcal{O}(10^{-3})$.
@@
Figure 2: The squared matrix elements of LO and $\lambda$ dependent EW corrections.
@@
The LO cross section of the VBF channel also exhibits a quadratic dependence on the trilinear Higgs coupling. The higher power dependence can be obtained by calculating the one-loop diagrams with an additional Higgs propagator. The calculation is standard except for the renormalization, which has been illustrated above. We have implemented our analytical results in the proVBFHH program [62, 41] and used the QCDLoop package [63] to evaluate the scalar one-loop integrals.
@@
3 Numerical results and updated constraints on Higgs boson self-coupling

### v2

@@
Improved constraints on Higgs boson self-couplings with quartic and cubic power dependencies in the cross section
Hai Tao Lia, Zong-Guo Sia, Jian Wanga,b111j.wang@sdu.edu.cn, Xiao Zhanga, Dan Zhaoa
aSchool of Physics, Shandong University, Jinan, Shandong 250100, China
bCenter for High Energy Physics, Peking University, Beijing 100871, China
The precise determination of the Higgs boson self-couplings is essential for understanding the mechanism behind electroweak symmetry breaking. However, due to the limited number of Higgs boson pair events at the LHC, only loose constraints have been established so far. Current constraints are based on the assumption that the cross section is a quadratic function of the trilinear Higgs self-coupling within the $\kappa$ framework. Incorporating higher-order quantum corrections from virtual Higgs bosons would significantly alter this function form, introducing new quartic and cubic power dependencies on the trilinear Higgs self-coupling. To derive this new function form, we propose a specialized renormalization procedure that tracks all Higgs self-couplings at each calculation step. Additionally, we introduce renormalization constants for the coupling modifiers within the $\kappa$ framework to ensure the cancellation of all ultraviolet divergences. With the new function forms of the cross sections in both the gluon-gluon fusion and vector boson fusion channels, the upper limit of $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ by the ATLAS (CMS) collaboration is reduced from 6.6 (6.49) to 5.4 (5.37). However, extracting a meaningful constraint on the quartic Higgs self-coupling $\lambda_{\rm 4H}$ from Higgs boson pair production data remains challenging. We also present the invariant mass distributions of the Higgs boson pair at different values of $\kappa_{\lambda}$, which could aid in setting optimal cuts for experimental analysis.
Keywords: Higgs boson self-couplings, higher power dependencies, precise predictions
@@ 1 Introduction
Following the discovery of the Higgs boson at the Large Hadron Collider (LHC) [1, 2], precise measurements of its properties, including mass, spin, and couplings to gauge bosons and fermions, have become critically important [3, 4, 5, 6, 7, 8, 9]. These measurements have thus far been consistent with the expectations of the Standard Model (SM) [10, 11]. However, the trilinear and quartic Higgs self-couplings, denoted as $\lambda_{\rm 3H}$ and $\lambda_{\rm 4H}$, respectively, which represent a fundamental aspect of the SM that connects the Higgs mechanism and the stability of our Universe [12], are still subject to large uncertainties.
@@
Significant efforts have been dedicated to improving the measurement of the Higgs self-coupling. The most direct approach involves measuring the cross section of Higgs boson pair production, predominantly through gluon-gluon fusion (ggF). At the leading order (LO), the process occurs via a top-quark loop. In the large top-quark mass ($m_{t}$) limit, the cross section of ggF Higgs boson pair production is known up to next-to-next-to-next-to-leading-order (N3LO) QCD corrections  [13, 14, 15, 16, 17], with the calculation of the soft gluon resummation effect also being studied [18, 19, 20]. When considering the full $m_{t}$ dependence, only next-to-leading-order (NLO) QCD corrections are available [21, 22, 23, 24], while estimates of the finite $m_{t}$ effects at next-to-next-to-leading-order (NNLO) have been conducted [25, 26, 27, 28, 29]. A comprehensive simulation of events necessitates a fully differential calculation of the Higgs boson pair production and decay to the $b\bar{b}\gamma\gamma$ final state at QCD NLO [30]. Furthermore, investigations into the NLO electroweak corrections have been explored [31, 32, 33, 34, 35, 36, 37]. The subdominant channel—vector boson fusion (VBF)—has also been computed up to N3LO in QCD [38, 39, 40, 41, 42].
@@
The current constraints on the trilinear Higgs self-coupling extracted from the Run 2 dataset of Higgs boson pair production at the LHC by the ATLAS and CMS experiments are $-0.6<\kappa_{\lambda_{\rm 3H}}<6.6$ [43] and $-1.24<\kappa_{\lambda_{\rm 3H}}<6.49$ [11], respectively, within the $\kappa$ framework [44], where $\kappa_{\lambda_{\rm 3H}}=\lambda_{\rm 3H}/\lambda_{\rm 3H}^{\rm SM}$ with $\lambda_{\rm 3H}^{\rm SM}$ being the SM value of the trilinear Higgs self-coupling. On the other hand, the process of single Higgs boson production and decay depends on the Higgs self-coupling only via higher-order electro-weak (EW) corrections and can also impose certain constraint [45, 46, 47, 48, 49, 50]. A measurement using the differential fiducial cross section in bins of the Higgs boson transverse momentum sets a constraint of $-5.4<\kappa_{\lambda_{\rm 3H}}<14.9$ [51]. Combined analyses of single- and double-Higgs production result in the constraint, $-0.4<\kappa_{\lambda_{\rm 3H}}<6.3$, at the $95\%$ confidence level, assuming that the new physics changes only the Higgs self-coupling [43].
@@
These constraints stem from considering the cross section of Higgs boson pair production as a function of the Higgs self-coupling. Indeed, even with higher-order QCD corrections, the cross section is a quadratic function of the trilinear Higgs self-coupling $\lambda_{\rm 3H}$. Nevertheless, higher-order EW corrections introduce contributions from Feynman diagrams containing one or more triple Higgs or quadruple Higgs vertices, leading to a distinct functional dependence on the Higgs self-coupling. Specifically, new quartic and cubic power dependencies on the trilinear Higgs self-coupling emerge, which has a significant impact on the constraints, given that the current upper limit is so large.
In practical calculations, maintaining an explicit dependence on the Higgs self-coupling can be challenging, as it is typically treated as a derived parameter in the conventional calculation of EW corrections, particularly during the renormalization process [52, 35]. Even if one can perform the renormalization by taking the Higgs self-coupling as a primary parameter, it is not clear how to implement the relation $m_{H}^{2}=2\lambda v^{2}$ and to rescale the Higgs self-coupling. Different choices lead to different expressions for the cross sections. Therefore, the calculation of the EW corrections in the SM can not be extended to the case with general $\lambda_{\rm 3H}$ and $\lambda_{\rm 4H}$, and the cross section with higher power (beyond quadratic) dependence on the Higgs self-couplings in the general $\kappa$ framework is still lacking.
To address this challenge, we propose a renormalization procedure that explicitly retains the Higgs self-couplings at each step and introduce renormalization of the coupling modifier. By combining it with the analytical and numerical calculations of the complex one-loop and two-loop amplitudes, respectively, we derive the cross sections of the Higgs boson pair production in both the ggF and VBF channels as functions of the Higgs self-couplings. Our findings indicate that incorporating higher power dependencies of the cross section on the Higgs self-couplings can reduce the upper limit of $\kappa_{\lambda_{\rm 3H}}$ by approximately 20%.
@@ 2 Renormalization in the $\kappa$ framework
The LO contribution to the ggF Higgs boson pair production $g(p_{1})g(p_{2})\to H(p_{3})H(p_{4})$ arises from the top-quark induced triangle and box Feynman diagrams, which are of order $\lambda_{\rm 3H}$ and $\lambda_{\rm 3H}^{0}$, respectively. Therefore, the LO cross section at the 13 TeV LHC can be expressed as
@@
Here, the subscript $\lambda_{\rm 3H}$ in $\kappa_{\lambda_{\rm 3H}}$ signifies the deviation of the trilinear Higgs self-coupling from its SM value. Below we will also introduce $\kappa_{\lambda_{\rm 4H}}$ to denote the modification of the quartic Higgs self-coupling. The SM cross section is recovered when $\kappa_{\lambda_{\rm 3H}}=1$ and $\kappa_{\lambda_{\rm 4H}}=1$. This quadratic functional form persists even with the inclusion of higher-order QCD corrections [53], e.g.,
$\sigma_{\rm ggF,NNLO-FT}^{\kappa_{\lambda}}=(~{}10.8~{}\kappa_{\lambda_{\rm 3H}}^{2}-49.6~{}\kappa_{\lambda_{\rm 3H}}+70.0~{})~{}~{}{\rm fb}.$ (2)
In this expression, the full one-loop real contributions are merged with other NNLO QCD corrections in the large $m_{t}$ limit. It is worth noting that the effects of QCD corrections are substantial, with each term in $\kappa_{\lambda_{\rm 3H}}$ more than doubling compared to the LO expression.
The EW corrections would change the above function form in two aspects. Firstly, the coefficients of the quadratic, linear, and constant terms are altered by the corrections induced by virtual gauge bosons. However, the overall impact is relatively minor, typically amounting to only a few percent, as reported in Ref. [35]. Given this negligible influence on the constraints related to the Higgs self-coupling, these corrections are deemed insignificant for the purposes of our study and are consequently omitted. Secondly, higher power or new dependence on the Higgs self-couplings arises due to the corrections induced by virtual Higgs bosons. One can see in Fig. 1 some typical two-loop Feynman diagrams, which give contributions of order $\lambda_{\rm 3H}^{3}$, $\lambda_{\rm 3H}^{2}$, $\lambda_{\rm 4H}\lambda_{\rm 3H}$, $\lambda_{\rm 4H}$ to the amplitudes. As a result, the cross section now contains new quartic and cubic powers of $\lambda_{\rm 3H}$ and starts to be sensitive to the quartic Higgs self-coupling $\lambda_{\rm 4H}$. It is the goal of our work to assess these corrections, denoted by $\delta\sigma_{\rm EW}^{\kappa_{\lambda}}$, and their impact on the constraints on the Higgs self-couplings.
@@
Our calculation of the two-loop diagrams proceeds as follows. We use FeynArts [54] to generate the Feynman diagrams and corresponding amplitudes. The amplitudes are written as a linear combination of two tensor structures with the coefficients called form factors [55]. After performing the Dirac algebra with FeynCalc [56, 57, 58], we are left with scalar integrals for each form factor. Rather than endeavoring to reduce all scalar integrals to master integrals and establishing differential equations for these master integrals, we opt to directly compute the scalar integrals for specific phase space points employing the numerical package AMFlow [59, 60]. This decision is motivated by the intricate nature of constructing differential equations, particularly with general kinematic dependencies, which can be exceedingly time-consuming. Even if the differential equation is obtained, the analytical solution seems not feasible with current technologies because all the propagators are massive. Numerical solutions of the differential equations often suffer from accuracy loss. By contrast, direct numerical calculation at each phase space point ensures accuracy. The primary challenge lies in covering the entire phase space efficiently. Fortunately, the process of $gg\to HH$ is dominated by the S-wave scattering, making the amplitude insensitive to the scattering angle. Moreover, its dependence on the scattering energy is also weak, except in very high-energy regions, as illustrated below. These characteristics allow us to generate a grid with a limited data set, which can be used to accurately calculate the amplitude at any point in the phase space.
Notice that the sum of all one-particle irreducible two-loop diagrams is finite. But the sum of one-particle reducible two-loop diagrams contains ultraviolet divergences. They will cancel after considering the contributions from the counter-terms in renormalization.
@@
where $Z_{v}$ is the renormalization constant for the vacuum expectation value. The renormalized Lagrangian in the $\kappa$ framework after EW gauge symmetry breaking is given by
$\mathcal{L}_{\rm H}^{\kappa}$ $=\frac{1}{2}Z_{\phi}(\partial_{\mu}H)^{2}-\left(-\frac{1}{2}Z_{\mu^{2}}Z_{\phi}Z_{v}^{2}\mu^{2}v^{2}+\frac{1}{4}Z_{\lambda}Z_{\phi}^{2}Z_{v}^{4}\lambda v^{4}\right)-(Z_{\lambda}Z_{\phi}^{2}Z_{v}^{3}\lambda v^{3}-Z_{\mu^{2}}Z_{\phi}Z_{v}\mu^{2}v)H$
@@
where the ellipsis represents the terms involving EW gauge bosons. Note that the letter $\lambda$ is solely used for the Higgs self-coupling in the SM but $\lambda_{\rm 3H}\equiv\kappa_{\lambda_{\rm 3H}}\lambda$ and $\lambda_{\rm 4H}\equiv\kappa_{\lambda_{\rm 4H}}\lambda$ denote the Higgs self-couplings that could be modified by new physics. We have added renormalization constants $Z_{\kappa_{\rm 3H}}$ and $Z_{\kappa_{\rm 4H}}$ for the coupling modifiers $\kappa_{\lambda_{\rm 3H}}$ and $\kappa_{\lambda_{\rm 4H}}$ to account for potential new physics effect in renormalization. Following the general principle of the $\kappa$ framework, we have assumed that the new physics does not affect the vacuum expectation value and Higgs mass when we rescale the Higgs self-couplings. The second term in the first line of Eq. (5) does not contain any field and thus can be safely dropped. Writing $Z=1+\delta Z$, the third term can be expanded as
@@
where we have neglected higher-order corrections that are products of two $\delta Z$’s. We choose the renormalization condition such that there is no tadpole contribution. This condition requires $\mu^{2}=\lambda v^{2}$ at the tree level, and $(\delta Z_{\mu^{2}}-\delta Z_{\lambda}-\delta Z_{\phi}-2\delta Z_{v})\mu^{2}v+T=0$ at the one-loop level with $T$ being the contribution from the one-loop tadpole diagrams. The vacuum expectation value appears always in the form of $Z_{\phi}^{1/2}Z_{v}v$, and is closely related to the massive gauge boson mass. Therefore $\delta Z_{v}$ would be determined only after considering the renormalization of the EW gauge sector. Since we focus on the corrections induced by the Higgs self-couplings, we can simply take $\delta Z_{v}+\delta Z_{\phi}/2=0$. We adopt dimensional regularization, i.e., setting the space-time dimension $d=4-2\epsilon$, to regulate the ultraviolet divergence and take $\mu_{R}$ as the renormalization scale. The tadpole diagram is evaluated to be
@@
$\frac{1}{2}(\partial_{\mu}H)^{2}-\mu^{2}H^{2}+\frac{1}{2}\delta Z_{\phi}(\partial_{\mu}H)^{2}-\left(\frac{3}{2}\delta Z_{\lambda}+\frac{5}{2}\delta Z_{\phi}-\frac{1}{2}\delta Z_{\mu^{2}}+3\delta Z_{v}\right)\mu^{2}H^{2}$
@@
where $s=(p_{1}+p_{2})^{2},\beta=\sqrt{1-4m_{H}^{2}/s}$ and $C_{0}[m_{H}^{2},m_{H}^{2},s,m_{H}^{2},m_{H}^{2},m_{H}^{2}]$ is a scalar integral that can be calculated by Package-X [61]. $\mathcal{M}_{gg\to H^{*}\to HH}^{\rm LO}$ is the LO amplitude which contains the Higgs self-coupling. In the SM, the divergences in the first line vanish and $\delta Z_{\kappa_{\rm 3H}}$ is not needed. In the general $\kappa$ framework, it is essential to include $\delta Z_{\kappa_{\rm 3H}}$ in renormalization. The presence of $\delta Z_{\kappa_{\rm 3H}}$ for general Higgs self-couplings demonstrates that the cross section in the $\kappa$ framework can not be derived from the result in the SM, as mentioned in the introduction. We adopt the $\overline{\rm MS}$ scheme to subtract the divergences. As a result, the coupling modifier is scale-dependent. If it is expressed in terms of the value at the scale $m_{H}$, then an additional contribution from its perturbative expansion cancels the above $\ln(\mu_{R}^{2}/m_{H}^{2})$ term exactly.
The $\kappa$ framework was initially established based on the signal strength obtained from experimental results. Here we have presented a field definition of the $\kappa$ framework in the Higgs sector so that one can perform higher-order calculations. A more systematical approach is to use the Higgs effective field theory (HEFT) with the electroweak chiral Lagrangian [62], which can be considered as an upgrade of the $\kappa$ framework to a quantum field theory and provides a general EFT description of the electroweak interactions with the presently known elementary particles under a cutoff scale around a few TeV [63]. The physical Higgs field $H$ is introduced as a singlet under $SU(2)_{L}\times U(1)_{Y}$ and the chiral symmetry. Our strategy, which includes the choice of unitary gauge and the implementation of the $\kappa$ parameter in the symmetry-broken phase, is equivalent to the application of HEFT in Higgs boson pair production. This equivalence can be easily verified by comparing our Lagrangian in eq. (5) and the one in eq. (2.5) of ref. [64].
@@
Now we are ready to compute the finite part of the squared amplitudes. We set the two-dimensional grid as a function of the Higgs velocity $\beta$ and $\cos\theta$ with $\theta$ the scattering angle. The range of $\beta$ is from 0 to 1, and $\cos\theta$ is also taken from 0 to 1 since the squared amplitudes are symmetric under $\theta\to\pi-\theta$. The grids for the LO squared amplitudes and ${\lambda}$ dependent EW corrections are illustrated in Fig. 2. We see that the squared amplitudes are stable against the change within the region $\beta<0.6$. For larger $\beta$, the LO squared amplitudes rise dramatically and then start to drop when $\beta>0.9$. By contrast, the $\lambda$ dependent correction first decreases and then increases when $\beta$ is larger than 0.85. These variations are mainly due to the large logarithms $\ln^{i}(1-\beta)$. And therefore we have constructed the grid as a function of $\ln^{i}(1-\beta)$ for $\beta\geq 0.96$. We have tested the grid by comparing the generated values [^2] The Lagrange interpolation method is used to generate the squared amplitudes at the phase space point not on the lattice. and the ones obtained by direct high-precision computation at some phase space points that are not on the grid lattice. We find good agreement at the per-mille level. We have used the grid to calculate the LO total cross section by performing the convolution with the parton distribution function (PDF) and phase space integrations. Comparing it to the result obtained using analytical expressions or the OpenLoops package [65, 66, 67], we find the relative difference lower than $\mathcal{O}(10^{-3})$.
@@
Figure 2: The LO squared amplitudes (left) and $\lambda$ dependent EW corrections (right).
@@
The LO cross section of the VBF channel also exhibits a quadratic dependence on the trilinear Higgs coupling. The higher power dependence can be obtained by calculating the one-loop diagrams with an additional Higgs propagator. The calculation is standard except for the renormalization, which has been elaborated above. We have implemented our analytical results in the proVBFHH program [68, 41] and used the QCDLoop package [69] to evaluate the scalar one-loop integrals.
@@
3 Numerical results and improved constraints on the Higgs boson self-coupling